# Mingle: Mixtures of Null-Space Gated Low-Rank Experts for Test-Time Continual Model Merging

### v1

@@ Abstract
Continual model merging integrates independently fine-tuned models sequentially without access to original training data, providing a scalable and efficient solution to continual learning. However, current methods still face critical challenges, notably parameter interference among tasks and limited adaptability to evolving test distributions. The former causes catastrophic forgetting of integrated tasks, while the latter hinders effective adaptation to new tasks. To address these, we propose Mingle, a novel framework for test-time continual model merging, which leverages test-time adaptation using a small set of unlabeled test samples from the current task to dynamically guide the merging process. Mingle employs a mixture-of-experts architecture composed of parameter-efficient, low-rank experts, enabling efficient adaptation and improving robustness to distribution shifts. To mitigate catastrophic forgetting, we propose Null-Space Constrained Gating, which restricts gating updates to subspaces orthogonal to prior task representations. This suppresses activations on old task inputs and preserves model behavior on past tasks. To further balance stability and adaptability, we design an Adaptive Relaxation Strategy, which dynamically adjusts the constraint strength based on interference signals captured during test-time adaptation. Extensive experiments on standard continual merging benchmarks demonstrate that Mingle achieves robust generalization, reduces forgetting significantly, and consistently surpasses previous state-of-the-art methods by 7-9% on average across diverse task orders.
@@ 1 Introduction
Continual learning seeks to incrementally adapt machine learning models to new tasks without forgetting previously learned knowledge, addressing the critical challenge of catastrophic forgetting [41]. However, conventional continual learning approaches typically require continuous access to original training data, raising significant concerns about privacy and substantial computational overhead due to retraining efforts, thus limiting their applicability in dynamic, data-sensitive environments.
@@
Despite its advantages in scalability, data privacy, and distributed training capabilities [42, 13, 55], existing CMM methods still encounter critical issues, notably severe parameter interference between tasks and limited adaptability to evolving test distributions. This parameter interference arises because, as fine-tuned models are incrementally merged, overlapping or conflicting parameter updates accumulate, resulting in severe forgetting of previously learned tasks. To mitigate this interference, recent methods introduce structural constraints such as orthogonal projection [68, 74], model linearization [37, 67], and pruning-based sparsification [82, 86]. However, their effectiveness diminishes as task count grows and interference becomes increasingly entangled. Moreover, models merged across tasks often fail to generalize effectively, particularly when facing unseen or shifting task conditions. These flaws result in severe forgetting of earlier tasks and substantial performance gaps compared to the upper bound achieved by individually fine-tuned models. As shown in Fig. 1, TA [23] suffers from large performance gaps and strong forgetting, reflected by low accuracy and negative backward transfer (BWT). OPCM [68] improves over TA via orthogonalized merging but still shows notable degradation.
Figure 1: Performance after eight task continual merging: accuracy on first four tasks and overall BWT.
@@
  • We formalize test-time continual model merging (TTCMM), a novel framework leveraging unlabeled test samples for merging independently fine-tuned models.
  • We propose Mingle, a TTCMM method with Adaptive Null-Space Constrained Gating to effectively balance stability and plasticity.
@@ 2 Related Work
Continual Learning.  Continual learning (CL) aims to mitigate catastrophic forgetting [41], where models forget prior tasks when learning new ones. Regularization-based methods constrain updates via importance weights [30, 88, 2, 29, 78], while distillation techniques preserve knowledge by aligning outputs or intermediate features [19, 11, 58, 50]. Replay-based methods store exemplars or generate surrogates using prompts, prototypes, or generators [53, 38, 75, 59, 51]. Dynamic architectures expand capacity through growth or ensembling [34, 94, 40]. Recent work integrates lightweight adapters or prompts into pre-trained models to enhance transfer with minimal updates [85, 22, 76]. CL has also been approached through model merging [77, 23, 27, 66]. One line remains close to conventional CL, incrementally fine-tuning models and merging them to reduce forgetting [39, 40, 15]. These methods typically require access to training data and involve sequential task-wise fine-tuning. In contrast, the continual model merging paradigm [27, 37, 5, 49, 68] merges independently fine-tuned models without revisiting prior data, offering greater scalability and privacy.
Model Merging.  Early work merged models via direct parameter averaging [70, 56], later refined by linear mode connectivity [12, 1]. Task Arithmetic (TA) [23] generalizes this by viewing models as task vectors to be summed. However, TA relies on weight disentanglement [47], which standard fine-tuning often violates, prompting structured training [27, 63]. Beyond averaging, interference-aware methods reweight or sparsify parameters [82, 86], or fuse models via distillation and clustering [83, 73]. LoRA-based tuning [20] introduces additional challenges due to parameter entanglement, spurring gradient-free or retrieval-based fusion strategies [21, 90, 91].
@@
Test-Time Adaptation.  TTA updates models at inference to counter test-time distribution shift. Early methods employed self-supervised objectives [64], entropy minimization [72], and regularized updates [61]. Online TTA adapts continuously [24], while batch-wise variants ignore temporal structure [16]. To improve stability, recent methods incorporate confidence filtering [46], EMA [10], partial updates [87], test-time augmentation [89], and adaptive BatchNorm [54]. In vision–language models, TTA leverages prompts or adapters [57, 14, 35]. Our approach draws on TTA to steer merging dynamics, aligning fused models with continuously evolving test distributions.
Discussion.   Most related to our method are MoE-Adapter [85] and WEMOE [66], both leveraging MoE architectures [26, 28]. MoE-Adapter follows conventional CL, where expert modules are embedded in the model and trained as a whole across tasks. In contrast, we adopt a model merging paradigm: experts are extracted from independently fine-tuned models and inserted without training. While WEMOE incorporates TTA, it is designed for multitask learning rather than continual scenarios, relying on simultaneous access to all models and data. By contrast, Mingle is explicitly tailored for the more challenging continual merging setup.
3 Mingle: Mixtures of Null-Space Gated Low-Rank Experts
@@ 3.1 Preliminaries
Existing Continual Merging Strategies.
Let $\theta_{0}$ denote the parameters of a pre-trained model. The corresponding task vector is defined as $\Delta\theta_{t}=\theta_{t}-\theta_{0}$.
@@ 3.2 Motivation and Theoretical Analysis
3.3 Low-Rank Expert Mixtures for Continual Model Merging
@@
Mixtures of Low-Rank Expert.  When a new task $t$ arrives, a dedicated expert $f_{t}$ and its gate $g_{t}$ are appended to the model. The output of a given $l$-th layer [^2] The layer index $l$ is omitted hereafter whenever it does not cause ambiguity. can be formulated as follows:
@@
where $u_{p}^{(t-1)}$ and $v_{q}^{(t-1)}$ are the $p$-th and $q$-th singular vectors from the singular value decomposition (SVD) of previous experts $\sum_{i=1}^{t-1}f_{i}(X)$. This projection removes previously learned directions to mitigate interference. We then apply a rank-$r$ truncated SVD for $\mathcal{P}^{(t-1)}(\Delta\theta_{t})$ to construct a low-rank expert [20]:
@@
Test-Time Adaptation.  Given class text embeddings $T=[T_{1},\dots,T_{|\mathcal{C}_{t}|}]\in\mathbb{R}^{|\mathcal{C}_{t}|\times d}$, where each $T_{i}$ is a prompt (e.g., “a photo of a dog”), the model computes logits via scaled cosine similarity: $p(y\mid x)=\mathrm{softmax}(z(x)),\ z(x)=\tau\cdot\mathrm{norm}(\phi_{\text{img}}(x))\cdot\mathrm{norm}(\phi_{\text{text}}(T))^{\top},$ where $\phi_{\text{img}}$ and $\phi_{\text{text}}$ are the image and text encoders, $\mathrm{norm}(\cdot)$ denotes $L_{2}$-normalization, and $\tau$ is a temperature scalar.
To encourage the merged model to retain task-specific behavior, we minimize the Kullback–Leibler divergence between its prediction and that of the corresponding individual fine-tuned model $\theta_{t}$. We define the adaptation objective as:
@@ 3.4 Adaptive Null-Space Constrained Gating for Interference Mitigation
Hard Null-Space Projection.  After finishing task $t$ we cache $l$-th layer activations on a seed buffer $\mathcal{D}_{t}^{\text{seed}}$ and estimate its covariance $C_{t}^{(l)}=\mathbb{E}[\theta_{t}^{\mathrm{merged},(l)}(X_{t})\,\theta_{t}^{\mathrm{merged},(l)}(X_{t})^{\top}]\in\mathbb{R}^{d\times d}$. Truncated SVD yields the top-$k$ dominant subspaces $\tilde{U}_{t}^{(l)}\in\mathbb{R}^{d\times k}$. Concatenating all previous tasks and orthonormalising: $U_{t}^{(l)}=\operatorname{orthonorm}\bigl{[}U_{t-1}^{(l)}|\tilde{U}_{t}^{(l)}\bigr{]}\in\mathbb{R}^{d\times tk}$, the hard projector is $P_{t}=I-U_{t}^{(l)}\,U_{t}^{(l)\top}\in\mathbb{R}^{d\times d}.$ To eliminate interference on task $\leq t-1$, the gate update for task $t$ is $W_{t}^{(g,l)}\leftarrow W_{t}^{(g,l)}-\eta\,P_{t-1}^{(l-1)}\nabla L_{t}^{(l)},$ which may also removes gradient components that are informative for task $t$ whenever its feature support overlaps $\operatorname{span}(U_{t-1})$.
@@
       $f_{t}=\text{SVD}~{}\!\bigl{(}\,\underbrace{\mathcal{P}^{(t-1)}(\theta_{t}-\theta_{0})}_{\text{use }\theta_{1}-\theta_{0}\text{ when }t=1}\bigr{)}=B_{t}A_{t}$
@@
       $\theta_{t}^{\text{merged}}=\theta_{t-1}^{\text{merged}}+g_{t}\!\cdot\!f_{t}$
      $\{W_{t}^{(g)},b_{t}^{(g)},S^{(0)}\}\leftarrow 0$
      for $m=1$ to total steps do
@@
                  $\nabla_{W_{t}^{(g)}}L\leftarrow\widetilde{P}_{t-1}\,\nabla_{W_{t}^{(g)}}L$
@@
            $W_{t}^{(g)}\leftarrow W_{t}^{(g)}-\eta\,\nabla_{W_{t}^{(g)}}L$
            $b_{t}^{(g)}\leftarrow b_{t}^{(g)}-\eta\,\nabla_{b_{t}^{(g)}}L$
@@
We maintain per-direction interference scores $S^{(m,l)}\in\mathbb{R}^{k}$ (initialized to 00) by applying an exponential moving average at each optimization step $m$:
@@
$W_{t}^{(g,l)}\;\leftarrow\;W_{t}^{(g,l)}-\eta\,\widetilde{P}_{t-1}^{(l-1)}\,\nabla L_{t}^{(l)}.$ (11)
Relaxing the projector inevitably allows more residual interference than the hard null-space variant, yet empirically the increase is minor and is offset by markedly higher plasticity (Tab. 4), indicating a favorable stability–plasticity balance. The overall procedure is outlined in Algo. 1.
@@ 4.1 Experimental Setup
Datasets and Models. Following [68], we evaluate our method on a diverse set of image classification tasks using CLIP-ViT models [52] as the backbone. Three task groups comprising 8, 14, and 20 tasks are considered to assess scalability. Specifically, we use ViT-B/32, ViT-B/16, and ViT-L/14 models, each fine-tuned on up to 20 downstream tasks. All resulting model checkpoints are publicly available via FusionBench [65]. To evaluate robustness to task order, we repeat experiments over 10 random permutations using seeds ranging from 42 to 51. To enable comparison with conventional continual learning methods, we adopt the Multi-domain Task-Incremental Learning (MTIL) benchmark [93], where eleven tasks are presented in alphabetical order.
@@
Implementation Details. We insert low-rank experts into the linear layers of the CLIP vision encoder. Two variants are considered: a full setup that modifies all attention and MLP linear layers, and a lightweight alternative applied only to attn.qkv and mlp.fc1. Unless stated otherwise, we adopt the full configuration. Each expert has rank $r=64$. The null-space constraint uses a subspace dimension $k=3$, with relaxation coefficients $\gamma=1$ and $\beta=0.99$. For each task, the merged model is adapted over 50 iterations using the Adam optimizer with a learning rate of 1e−4 and a batch size of 16. We draw 5 unlabeled test data per class, with no access to data from previous tasks.

### v2

# Mingle: Mixture of Null-Space Gated Low-Rank Experts for Test-Time Continual Model Merging

@@ Abstract
Continual model merging integrates independently fine-tuned models sequentially without access to the original training data, offering a scalable and efficient solution for continual learning. However, existing methods face two critical challenges: parameter interference among tasks, which leads to catastrophic forgetting, and limited adaptability to evolving test distributions. To address these issues, we introduce the task of Test-Time Continual Model Merging (TTCMM), which leverages a small set of unlabeled test samples during inference to alleviate parameter conflicts and handle distribution shifts. We propose Mingle, a novel framework for TTCMM. Mingle employs a mixture-of-experts architecture with parameter-efficient, low-rank experts, which enhances adaptability to evolving test distributions while dynamically merging models to mitigate conflicts. To further reduce forgetting, we propose Null-Space Constrained Gating, which restricts gating updates to subspaces orthogonal to prior task representations, thereby suppressing activations on old tasks and preserving past knowledge. We further introduce an Adaptive Relaxation Strategy that adjusts constraint strength dynamically based on interference signals observed during test-time adaptation, striking a balance between stability and adaptability. Extensive experiments on standard continual merging benchmarks demonstrate that Mingle achieves robust generalization, significantly reduces forgetting, and consistently surpasses previous state-of-the-art methods by 7–9% on average across diverse task orders. Our code is available at: https://github.com/zihuanqiu/MINGLE
@@ 1 Introduction
Continual learning aims to incrementally adapt machine learning models to new tasks without forgetting previously learned knowledge, addressing the critical challenge of catastrophic forgetting [43]. However, conventional continual learning approaches typically require continuous access to original training data, raising significant concerns about privacy and substantial computational overhead due to retraining efforts, thus limiting their applicability in dynamic, data-sensitive environments.
@@
Despite its advantages in scalability, data privacy, and distributed training capabilities [44, 14, 57], existing CMM methods still encounter critical issues, notably severe parameter interference between tasks and limited adaptability to evolving test distributions. This parameter interference arises because, as fine-tuned models are incrementally merged, overlapping or conflicting parameter updates accumulate, resulting in severe forgetting of previously learned tasks. To mitigate this interference, recent methods introduce structural constraints such as orthogonal projection [70, 78], model linearization [38, 69], and pruning-based sparsification [86, 90]. However, their effectiveness diminishes as task count grows and interference becomes increasingly entangled. Moreover, models merged across tasks often fail to generalize effectively, particularly when facing unseen or shifting task conditions. These flaws result in severe forgetting of earlier tasks and substantial performance gaps compared to the upper bound achieved by individually fine-tuned models. As shown in Fig. 1, TA [24] suffers from large performance gaps and strong forgetting, reflected by low accuracy and negative backward transfer. OPCM [70] improves over TA via orthogonalized merging but still shows notable degradation.
Figure 1: After 8-task continual merging: accuracy on first four tasks and overall BWT.
@@
  • We formalize test-time continual model merging (TTCMM), a novel task that leverages unlabeled test samples to merge independently fine-tuned models.
  • We propose Mingle, a TTCMM framework with Adaptive Null-Space Constrained Gating to effectively balance stability and plasticity.
@@ 2 Related Work
Continual Learning. Continual learning (CL) seeks to mitigate catastrophic forgetting [43], where learning new tasks overwrites prior knowledge. Regularization-based methods constrain updates with importance weights [31, 92, 2, 30, 82], while distillation aligns outputs to preserve knowledge [20, 12, 60, 52]. Replay methods store exemplars or generate surrogates with prompts, prototypes, or generators [55, 39, 79, 61, 53], and dynamic architectures expand capacity via growth or ensembling [35, 98, 42]. Recent work leverages lightweight adapters or prompts in pre-trained models for efficient transfer [89, 23, 80]. Model merging offers an alternative route. Some methods remain close to conventional CL by sequentially fine-tuning and merging models to reduce forgetting [41, 42, 16], typically requiring training data. In contrast, continual model merging [28, 38, 5, 51, 70] merges independently fine-tuned models without revisiting prior data, enabling greater scalability and privacy.
Model Merging. Early work merged models via direct parameter averaging [72, 58], later refined by linear mode connectivity [13, 1]. Wortsman et al. [81] showed that weight averaging can also enhance robustness and out-of-distribution generalization. Task Arithmetic (TA) [24] views models as task vectors to be summed, but relies on weight disentanglement [49], often violated under standard fine-tuning, motivating structured training [28, 65]. Beyond averaging, interference-aware methods reweight or sparsify parameters [86, 90], or fuse models via distillation and clustering [88, 76]. LoRA-based tuning [21] introduces additional entanglement challenges, spurring gradient-free or retrieval-based strategies [22, 94, 95]. More recently, dynamic merging [68, 40] adapts parameters conditioned on inputs, achieving higher flexibility and performance, but remains limited to multi-task fusion and unexplored in continual settings.
@@
Test-Time Adaptation. TTA adapts models at inference to mitigate distribution shift. Early approaches used self-supervised objectives [66], entropy minimization [75], or regularized updates [63]. Online TTA adapts continuously [25], while batch-wise variants ignore temporal structure [17]. To enhance stability, later work introduced confidence filtering [48], EMA [11], partial updates [91], test-time augmentation [93], and adaptive BatchNorm [56]. For vision–language models, TTA often employs prompts or adapters [59, 15, 36]. We draw on TTA to guide merging, aligning fused models with evolving test distributions.
Relation to Prior Work. Most related to our work are MoE-Adapter [89] and WEMOE [68], both built on MoE architectures [27, 29]. MoE-Adapter follows conventional continual learning, embedding expert modules that are jointly trained across tasks. In contrast, we adopt a model-merging paradigm, where experts are extracted from independently fine-tuned models and inserted without further training. WEMOE incorporates test-time adaptation but targets multitask learning, assuming simultaneous access to all models and data. By contrast, Mingle is tailored for the more challenging continual merging setting.
3 Mingle: Mixture of Null-Space Gated Low-Rank Experts
@@ 3.1 Preliminaries
Existing Continual Merging Strategies. Let $\theta_{0}$ denote the parameters of a pre-trained model. The corresponding task vector is defined as $\Delta\theta_{t}=\theta_{t}-\theta_{0}$.
@@ 3.2 Motivation and Theoretical Analysis
3.3 Low-Rank Expert Mixture for Continual Model Merging
@@
Mixture of Low-Rank Expert. When a new task $t$ arrives, a dedicated expert $f_{t}$ and its gate $g_{t}$ are appended to the model. The output of a given $l$-th layer [^2] The layer index $l$ is omitted hereafter whenever it does not cause ambiguity. can be formulated as follows:
@@
where $u_{p}^{(t-1)}$ and $v_{q}^{(t-1)}$ are the $p$-th and $q$-th singular vectors from the singular value decomposition (SVD) of previous experts $\sum_{i=1}^{t-1}f_{i}(X)$, and $\alpha$ denotes the effective rank of previous experts. This projection removes previously learned directions to mitigate interference. We then apply a rank-$r$ truncated SVD for $\mathcal{P}^{(t-1)}(\Delta\theta_{t})$ to construct a low-rank expert [21]:
@@
Test-Time Adaptation. To encourage the merged model to retain task-specific behavior, we minimize the Kullback–Leibler divergence between its prediction and that of the corresponding individual fine-tuned model $\theta_{t}$. We define the adaptation objective as:
@@ 3.4 Adaptive Null-Space Constrained Gating for Interference Mitigation
Hard Null-Space Projection. After completing task $t$, we cache the $l$-th layer inputs in a seed buffer $\mathcal{D}_{t}^{\text{seed}}$ and estimate their covariance $\mathrm{Cov}_{t}^{(l)}\in\mathbb{R}^{d\times d}$. Applying truncated SVD yields the top-$k$ dominant subspaces $\tilde{U}_{t}^{(l)}\in\mathbb{R}^{d\times k}$. We then concatenate these with the subspaces from all previous tasks and orthonormalize: $U_{t}^{(l)}=\operatorname{orthonorm}\bigl[U_{t-1}^{(l)}|\tilde{U}_{t}^{(l)}\bigr]\in\mathbb{R}^{d\times tk}$. The hard projector is $P_{t}=I-U_{t}^{(l)}\,U_{t}^{(l)\top}\in\mathbb{R}^{d\times d}.$ To suppress interference from tasks $\leq t-1$, the gating update for task $t$ is $W_{t}^{(g,l)}\leftarrow W_{t}^{(g,l)}-\eta\,\nabla L_{t}^{(l)}P_{t-1}^{(l)},$. However, this projection may also discard gradient components that are informative for task $t$ whenever its feature support overlaps with $\operatorname{span}(U_{t-1})$.
@@
 $f_{t}={\small\textsc{SVD}_{\textsc{trunc}.}}~\!\bigl(\,\underbrace{\mathcal{P}^{(t-1)}(\theta_{t}-\theta_{0})}_{\scriptsize\text{use }\theta_{1}-\theta_{0}\text{ when }t=1}\bigr)=B_{t}A_{t}$
@@
 $\theta_{t}^{\text{merged}}=\theta_{t-1}^{\text{merged}}+g_{t}\!\cdot\!f_{t},\;\{W_{t}^{(g)},b_{t}^{(g)},S^{(0)}\}\leftarrow 0$
 for $m=1$ to total iterations do
@@
       $\nabla_{W_{t}^{(g)}}L_{t}\leftarrow\nabla_{W_{t}^{(g)}}L_{t}\,\widetilde{P}_{t-1}$
@@
    $W_{t}^{(g)}\leftarrow W_{t}^{(g)}-\eta\,\nabla_{W_{t}^{(g)}}L_{t}$
    $b_{t}^{(g)}\leftarrow b_{t}^{(g)}-\eta\,\nabla_{b_{t}^{(g)}}L_{t}$
@@
We maintain per-direction interference scores $S^{(m,l)}\in\mathbb{R}^{k}$ (initialized to 0) by applying an exponential moving average at each iteration $m$:
@@
$W_{t}^{(g,l)}\;\leftarrow\;W_{t}^{(g,l)}-\eta\,\nabla L_{t}^{(l)}\,\widetilde{P}_{t-1}^{(l)}.$ (11)
Relaxing the projector inevitably allows more residual interference than the hard null‐space variant, yet empirically the increase is minor and is offset by markedly higher plasticity (Tab. 5), indicating a favorable stability–plasticity balance. The overall procedure is outlined in Algo. 1.
@@ 4.1 Experimental Setup
Datasets and Models.  Following [70], we evaluate on image-classification tasks with CLIP-ViT backbones [54]. We consider 8, 14, and 20-task groups using ViT-B/32, ViT-B/16, and ViT-L/14 models, each fine-tuned on up to 20 downstream tasks, with checkpoints from FusionBench [67]. To assess order sensitivity, we repeat experiments over 10 random seeds (42–51). For comparison with conventional CL, we use the Multi-domain Task-Incremental Learning (MTIL) benchmark [97] with eleven vision tasks. Beyond vision, we evaluate on eight GLUE language tasks [74] with a Flan-T5-base backbone [6].
@@
Implementation Details.  We insert low-rank experts into the CLIP vision encoder. Two variants are used: a full setup modifying all attention and MLP layers, and a lightweight one on attn.qkv and mlp.fc1. All experiments share a single set of global hyper-parameters across models and task orders. Each expert has rank $r=64$; the null-space constraint uses $k=3$, $\gamma=1$, and $\beta=0.99$. Adaptation runs for 50 iterations with Adam (lr $1\text{e-4}$, batch size 16). For vision tasks we use 5 unlabeled samples per class, and for NLP tasks 100 in total, all without access to prior-task data.